A class of 125 students from civil, mechanical and electrical engineering backgrounds is divided into 2 sections (sections A and B). Section A contains of 18 civil engineers and a total of 50 students. Section B contains 44 mechanical engineers. There are a total 40 civil engineers and 60 mechanical engineers in the class.

What is the probability that a random student is a Civil engineer and from section B? Please enter the answer to 3 decimal places.
Answer #1

Total number of students = 125

Total number of civil engineers = 40

Number of civil engineers in Section A = 18

Number of civil engineers in Section B = 40 - 18 = 22

P(civil engineer and from Section B) = 22/125

= 0.176